OpenStudy (hanna14):

Please I will medal and fan!!!!!!!!!! I am soooo desperate. The coordinates of the vertices of triangle ABC are A(-1,3), B(1,2) and C(-3,-1). Determine the slope of each side of the triangle and use that information to determine if the triangle is a right triangle or not.

OpenStudy (sooobored):

a right triangle is a triangle with one right angle in it a right angle is a 90 degree angle 2 perpendicular lines will create a 90 degree angle in order for 2 lines to be perpendicular to each other, their slopes must be negative reciprocals to each other the slope between two point cans be given as \[m= \frac{y_2 - y_1}{x_2-x_1}\]

OpenStudy (sooobored):

you have 3 points, A, B, C you can create 3 line segments to create the triangle AB, BC, AC what is the slope of AB? what is the slope of BC? what is the slope of AC?

OpenStudy (hanna14):

Ok I am trying to find the slopes now

OpenStudy (hanna14):

1) Slope of AB is =1/2? 2) Slope of BC is 3/4? 3) Slope of AC is 2?

OpenStudy (hanna14):

@sooobored

OpenStudy (sooobored):

ok they look fine, are any of the slopes negative reciprocals to each other?

OpenStudy (sooobored):

negative reciprocal of a rational number \[\frac{a}{b} \rightarrow -\frac{b}{a}\]

OpenStudy (sooobored):

 sorry, slope of AB is wrong

OpenStudy (sooobored):

forgot a negative sign in my head

OpenStudy (sooobored):

unless that equal sign "=" is suppose to be a negative sign "-" cause slope of AB should be -1/2

OpenStudy (hanna14):

Oh yeah that is supposed to be negative thanks!

OpenStudy (sooobored):

well to finish off the question 2 is a negative reciprocal to -1/2 thus the line segment AB and AC just be perpendicular to each other and thus meaning that the triangle is a right triangle since one of its sides is a right angle/90 degrees/ perpendicular